# Latent Plan Transformer for Trajectory Abstraction: Planning as Latent Space Inference

### v3

@@ 1 Introduction
Decision Transformer (DT) (Chen et al., 2021) and some concurrent work (Janner et al., 2021) have popularized the research agenda of decision-making via generative modeling. The general idea is to consider decision-making as a generative process that takes in a representation of the task objective, e.g. the rewards or returns of a trajectory, and outputs a representation of the trajectory. Intuitively, a purposeful decision-making process should shift the trajectory distribution towards regimes with higher returns. In the classical decision-making literature, this is achieved by two interweaving processes, policy evaluation and policy improvement (Sutton and Barto, 2018). Policy evaluation promotes consistency in the estimated correlations between the trajectories and the returns. In DT, this is realized by the maximum likelihood estimation (MLE) of the joint distribution of sequences consisting of states, actions, and return-to-gos (RTG). Policy improvement shifts the distribution to improve the status quo expectation of the returns. In DT, this is naturally entailed since the policy is a distribution of actions conditioned on step-wise RTGs.
In this work, we are interested in the problem of planning. Among various ways to identify planning as a special class of decision-making problems, we pay particular attention to its data specification and inductive biases. As designing step-wise rewards requires significant effort and domain expertise, we focus on the problem of learning from trajectory-return pairs, where a trajectory is a sequence of states and actions, and the return is its total rewards. This design choice forces the agents to predict into the long-term future and figure out step-wise credits by themselves. A competitive Temporal Difference (TD) learning baseline, CQL (Kumar et al., 2020), was reported to be fragile under this data specification (Chen et al., 2021).
@@
In this work, we introduce the Latent Plan Transformer (LPT), a novel generative model featuring a latent vector modeled by a neural transformation of Gaussian white noise, a Transformer-based policy conditioned on this latent vector and a return estimation model. LPT is learned by maximum likelihood estimation (MLE). Given an expected return, posterior inference of the latent vector in LPT is an explicit process for iterative refinement of the plan. The inferred latent variable replaces RTG in the conditioning of the auto-regressive policy, providing richer information about the anticipated future. We further develop a mode-seeking sampling scheme that strongly enforce the temporal consistency for long-range planning, which is particularly effective in stitch trajectory, i.e., to compose parts of sub-optimal trajectories to reach far beyond (Fu et al., 2020). LPT demonstrates competitive performance in Gym-Mujoco locomotion, Franka kitchen, goal-reaching tasks in maze2d and antmaze, and a contingent planning task Connect Four. These empirical results support that latent variable inference can enable and improve planning in the absence of step-wise rewards.
@@ 2 Background
However, $RTG$ becomes less reliable when rewards are sparse or have non-trivial relations with the return. Distributing the return to each step is a credit assignment problem. Consider an example of an ideal credit assignment mechanism: When students receive partial credits for their incomplete answers, it’s more fair to give points equal to the full marks minus the expected points for all possible ways to finish the answer, rather than assuming students have no knowledge of the remaining parts. This credit assignment mechanism can be formalized as, $RTG_{t}^{Q}=\sum_{k=t}^{K}r(s_{\leq k},a_{\leq k})+\mathbb{E}[Q(s_{\leq K},a_{\leq K})]$. Here $Q$ can be estimated using deep TD learning with multi-step returns. Yamagata et al. (2023) instantiate a Markovian version and demonstrate improvement in trajectory stiching.
@@ 3.1 Model
$z$$z$$z_{0}$$z$$\tau$$z$$y$$z=U_{\alpha}{(z_{0})}$$p_{\alpha}(z)$$p_{\beta}(\tau|z)$$p_{\gamma}(y|z)$Cross-attentionCausal Transformer×$\times N$$z$$\tau=(s_{1},a_{1},s_{2},a_{2},\dots,s_{H},a_{H})$$a_{1},a_{2},\dots,\cdots,a_{H}$
@@
[…]
@@
The prior model $p_{\alpha}(z)$ is an implicit generator, defined as a learnable neural transformation of an isotropic Gaussian, $z=U_{\alpha}(z_{0})$ and $z_{0}\sim\mathcal{N}(0,I_{d})$. $U_{\alpha}(\cdot)$ is an expressive neural network, such as the UNet (Ronneberger et al., 2015). This approach is inspired by, yet contrasts with Pang et al. (2020), wherein the latent space prior is modeled as an Energy-based Model (EBM) (Xie et al., 2016). While EBM offers explicit unnormalized density, its sampling process is complex. Conversely, our model provides an implicit density with simpler sampling.
The trajectory generator $p_{\beta}(\tau|z)$ is a conditional autoregressive model with finite context $K$, $p_{\beta}(\tau|z)=\prod_{t=1}^{H}p_{\beta}(\tau_{(t)}|\tau_{(t-K)},...,\tau_{(t-1)},z)$ where $\tau_{(t)}=(s_{t},a_{t})$. It can be parameterized by a causal Transformer with parameter $\beta$, similar to Decision Transformer (Chen et al., 2021). Specifically, the latent variable $z$ is included in trajectory generation using cross-attention, as shown in Fig. 1 and controls each step of the autoregressive trajectory generation as $p_{\beta}(a_{t}|s_{t-K:t},a_{t-K:t-1},z)$. The action is assumed to follow a single-mode Gaussian distribution, i.e. $a_{t}\sim\mathcal{N}(g_{\beta}(s_{t-K:t},a_{t-K:t-1},z),I_{|A|})$.
@@ 3.2 Offline Learning
The full derivation of the learning method is in Section A.1. Let $\delta_{\alpha},\delta_{\beta},\delta_{\gamma}$ represent the expected gradients of $L(\theta)$ with respect to the model parameters $\alpha,\beta,\gamma$, respectively. The learning gradients for each component are formulated as follows.
@@
Estimating these expectations requires Markov Chain Monte Carlo (MCMC) sampling of the posterior distribution $p_{\theta}(z_{0}|\tau,y)$. We use the Langevin dynamics (Neal, 2011) for MCMC sampling, iterating as follows for a target distribution $\pi(z)$:
@@
The sampling process starts by initializing $z_{0}^{k=0}$ from a standard normal distribution ${\mathcal{N}}(0,I_{d})$. We then apply $N$ steps of Langevin dynamics (e.g., $N=15$) to approximate the posterior distribution, making our learning algorithm an approximate MLE. For a theoretical understanding of this noise-initialized finite-step MCMC, see Pang et al. (2020); Nijkamp et al. (2020); Xie et al. (2023). However, for large horizons (e.g.,$H$=1000), this method becomes slow and memory-intensive. To mitigate this, we adopt the persistent Markov Chain (PMC) (Tieleman, 2008; Xie et al., 2016; Han et al., 2017), which amortizes sampling across training iterations. During training, $z_{0}^{k=0}$ is initialized from the previous iteration and the number of updates is reduced to $N=2$ steps. See Section A.2 for training and architecture details.
@@ 3.3 Planning as Inference
This posterior sampling is achieved using Langevin dynamics similar to the training process. Specifically, we replace the target distribution in Eq. 5 with $p_{\theta}(z_{0}|y)$ and run MCMC for a fixed number of steps. Sampling from $p_{\theta}(z_{0}|y)$ eliminates the need for expensive back-propagation through the trajectory generator $p_{\beta}(\tau|z)$.
@@
Inspired by the classifier guidance (CG) (Dhariwal and Nichol, 2021; Ho and Salimans, 2022) in conditional diffusion models, we introduce a guidance weight $w$ to the original posterior in Eq. 6
@@
which has the score $\nabla_{z_{0}}\log\tilde{p}_{\theta}(z_{0}|y)=\nabla_{z_{0}}{\log p_{0}(z_{0})}+w\nabla_{z_{0}}{\log p_{\gamma}(y|z)}$. This guidance weight $w$ controls the interpolation between exploration and exploitation. When $w=1$, the sampled plans collectively represent the posterior density and account for Bayesian uncertainty, resulting in a provably efficient exploration scheme (Osband and Van Roy, 2017). When $w>1$, the sampled plans are more concentrated around the modes of the posterior distribution, which are plans more likely to the agent. The larger the value of $w$, the more confident the agent becomes, and the stronger the inclination towards exploitation.
@@
  Input: Learning iterations $T$, initial parameters $\theta_{0}=(\alpha_{0},\beta_{0},\gamma_{0})$, offline training samples $\mathcal{D}=\{\tau_{i},y_{i}\}_{i=1}^{n}$, posterior sampling step size $s$, the number of steps $N$, and the learning rate $\eta_{0},\eta_{1},\eta_{2}$.
@@
     1.Posterior sampling: For each $(\tau_{i},y_{i})$, sample $z_{0}\sim{p}_{\theta_{t}}(z_{0}|\tau_{i},y_{i})$ using Eq. 5 with $N$ steps and step-size $s$, where the target distribution $\pi$ is ${p}_{\theta_{t}}(z_{0}|\tau_{i},y_{i})$.
     2.Learn prior model $p_{\alpha}(z)$, trajectory generator $p_{\beta}(\tau|z)$ and return predictor $p_{\gamma}(y|z)$: $\alpha_{t+1}=\alpha_{t}+\eta_{0}\frac{1}{n}\sum_{i}\delta_{\alpha}(\tau_{i},y_{i})$; $\beta_{t+1}=\beta_{t}+\eta_{1}\frac{1}{n}\sum_{i}\delta_{\beta}(\tau_{i},y_{i})$; $\gamma_{t+1}=\gamma_{t}+\eta_{2}\frac{1}{n}\sum_{i}\delta_{\gamma}(\tau_{i},y_{i})$ as in Section 3.2.
@@
  Input: Expected return $y$, a trained model on offline dataset $\theta$, posterior sampling step size $s$ and the number of steps $N$, Horizon $H$ and an evaluation environment.
@@
     Sample $z_{0}\sim\tilde{p}_{\theta}(z_{0}|y)$ as in Eq. 7 using Eq. 5 with $N$ steps and step size $s$, where the target distribution $\pi$ is replaced by $\tilde{p}_{\theta}(z_{0}|y)\propto p_{0}(z_{0})p_{\gamma}(y|z=U_{\alpha}(z_{0}))^{w}$ and $z=U_{\alpha}(z_{0})$.
@@
     Sample $z_{0}\sim p_{\theta}(z_{0}|y)$ as in Eq. 6 using Eq. 5 with $N$ steps and step size $s$, where $\pi$ is replaced by $p_{\theta}(z_{0}|y)\propto p_{0}(z_{0})p_{\gamma}(y|z=U_{\alpha}(z_{0}))$ and $z=U_{\alpha}(z_{0})$.
@@ 4 A Sequential Decision-Making Perspective
However, behavior cloning is believed to suffer from drifting errors since it ignores covariate shifts in future steps (Ross and Bagnell, 2010). This concern is unique to sequential decision-making, as the agent cannot control the next state from a stochastic environment, like generating the next text token.
@@
Our modeling of the latent trajectory variable $z$ provides an alternative solution to the temporal consistency issue. Eq. 4 is minimizing the KL divergence
@@
where $p_{\bar{\theta}}^{y}(z|\tau)={p^{y}_{\mathcal{D}}(\tau,z)}/{p^{y}_{\mathcal{D}}(\tau)}$ and $\bar{\theta}=\theta$ highlights these distributions have the same parameterization as $p_{{\theta}}^{y}$ but are wrapped with stop_grad() operator when calculating gradients for $\theta$ (Han et al., 2017). Comparing Eqs. 8 and 9, it is now clear that $z$ plays a similar role as $RTG$ in promoting temporal consistency in autoregressive models. Uniquely, $p_{\bar{\theta}}^{y}(z|\tau)$ is the temporal abstraction intrinsic to the model, in contrast to step-wise rewards. From a sequential decision-making perspective, $z$ is effectively a plan that the agent is persistent to. From a generative modeling perspective, $z$ from different trajectory modes would decompose the density $p^{y}(a_{t}|s_{0:t},a_{0:t-1})$, relieving the burden of learning the autoregressive policy $p_{\beta}(a_{t}|s_{0:t},a_{0:t-1},z)$.
@@
One caveat is that the transition model estimation should not be conditioned on $y$. Mixing up more trajectory regimes could provide additional regularization for its estimation and generalization. Actually, environment stochasticity is a more concerning issue for autoregressive behavior cloning, as highlighted by Yang et al. (2022); Paster et al. (2022); Štrupl et al. (2022); Brandfonbrener et al. (2022); Villaflor et al. (2022); Eysenbach et al. (2022). Among them, Yang et al. (2022) pinpoints the issue by viewing $RTG$s as deterministic latent trajectory variables, closely related to what we present here. Uniquely, the latent variable $z$ in our model is inherently multi-modal (hence very non-deterministic) and ignorant of step-wise rewards. We postulate that the overfitting issue might be mitigated. This is validated by our empirical study inspired by Paster et al. (2022).
@@
Although RTG prediction and plan prediction both promote temporal consistency, they function very differently when mixing trajectories from multiple return-conditioned regimes. RTG prediction is a supervised learning over the joint distribution $p_{\mathcal{D}}(\tau,\rho)$. Simply mixing trajectories from multiple regimes can’t encourage generalization to trajectories that are stitched with those in the dataset. Yamagata et al. (2023) propose to resolve this by replacing $RTG$ with $RTG^{Q}$. Intuitively, this augments the distribution $p_{\mathcal{D}}(\tau,\rho)$ with $p_{\mathcal{D}}(\tau^{\prime},\rho^{Q})$, where $\tau^{\prime}$ denotes trajectories covered by the offline dynamic programming, such as Q learning, and $\rho^{Q}=(RTG_{0}^{Q},RTG_{1}^{Q},...Q_{H})$. It significantly improves tasks requiring trajectory stitching. Conversely, plan prediction is an unsupervised learning as it samples from $p_{\mathcal{D}}(\tau,y)p_{\bar{\theta}}(z|\tau,y)$. As $z$ contains more trajectory-related information than step-wise $RTG$s, trajectories lying outside of $p_{\mathcal{D}}(\tau,\rho)$ may be in-distribution for $p_{\mathcal{D}}(\tau,y)p_{\bar{\theta}}(z|\tau,y)$. The return prediction training further shapes the representation of $z$, which can be benefited from denser coverage of $y$. With more return values covered, we may count on neural networks’ strong interpolation capability to shift the trajectory distribution with $y$-conditioning.

### v4

@@ 1 Introduction
Decision Transformer (DT) [Chen et al., 2021] and some concurrent work [Janner et al., 2021] have popularized the research agenda of decision-making via generative modeling. The general idea is to consider decision-making as a generative process that takes in a representation of the task objective, e.g. the rewards or returns of a trajectory, and outputs a representation of the trajectory. Intuitively, a purposeful decision-making process should shift the trajectory distribution towards regimes with higher returns. In the classical decision-making literature, this is achieved by two interweaving processes, policy evaluation and policy improvement [Sutton and Barto, 2018]. Policy evaluation promotes consistency in the estimated correlations between the trajectories and the returns. In DT, this is realized by the maximum likelihood estimation (MLE) of the joint distribution of sequences consisting of states, actions, and return-to-gos (RTG). Policy improvement shifts the distribution to improve the status quo expectation of the returns. In DT, this is naturally entailed since the policy is a distribution of actions conditioned on step-wise RTGs.
In this work, we are interested in the problem of planning. Among various ways to identify planning as a special class of decision-making problems, we pay particular attention to its data specification and inductive biases. As designing step-wise rewards requires significant effort and domain expertise, we focus on the problem of learning from trajectory-return pairs, where a trajectory is a sequence of states and actions, and the return is its total rewards. This design choice forces the agents to predict into the long-term future and figure out step-wise credits by themselves. A competitive Temporal Difference (TD) learning baseline, CQL [Kumar et al., 2020], was reported to be fragile under this data specification [Chen et al., 2021].
@@
In this work, we introduce the Latent Plan Transformer (LPT), a novel generative model featuring a latent vector modeled by a neural transformation of Gaussian white noise, a Transformer-based policy conditioned on this latent vector and a return estimation model. LPT is learned by maximum likelihood estimation (MLE). Given an expected return, posterior inference of the latent vector in LPT is an explicit process for iterative refinement of the plan. The inferred latent variable replaces RTG in the conditioning of the auto-regressive policy, providing richer information about the anticipated future. We further develop a mode-seeking sampling scheme that strongly enforce the temporal consistency for long-range planning, which is particularly effective in stitch trajectory, i.e., to compose parts of sub-optimal trajectories to reach far beyond [Fu et al., 2020]. LPT demonstrates competitive performance in Gym-Mujoco locomotion, Franka kitchen, goal-reaching tasks in maze2d and antmaze, and a contingent planning task Connect Four. These empirical results support that latent variable inference can enable and improve planning in the absence of step-wise rewards.
@@ 2 Background
However, $RTG$ becomes less reliable when rewards are sparse or have non-trivial relations with the return. Distributing the return to each step is a credit assignment problem. Consider an example of an ideal credit assignment mechanism: When students receive partial credits for their incomplete answers, it’s more fair to give points equal to the full marks minus the expected points for all possible ways to finish the answer, rather than assuming students have no knowledge of the remaining parts. This credit assignment mechanism can be formalized as, $RTG_{t}^{Q}=\sum_{k=t}^{K}r(s_{\leq k},a_{\leq k})+\mathbb{E}[Q(s_{\leq K},a_{\leq K})]$. Here $Q$ can be estimated using deep TD learning with multi-step returns. Yamagata et al. [2023] instantiate a Markovian version and demonstrate improvement in trajectory stiching.
@@ 3.1 Model
[…]
@@
where $\theta=(\alpha,\beta,\gamma)$. LPT approximates the data distribution $p_{\mathrm{data}}(\tau,y)$ using the marginal distribution $p_{\theta}(\tau,y)=\int p_{\theta}(\tau,y,z)dz$. It also establishes a generation process,
@@
The prior model $p_{\alpha}(z)$ is an implicit generator, defined as a learnable neural transformation of an isotropic Gaussian, $z=U_{\alpha}(z_{0})$ and $z_{0}\sim\mathcal{N}(0,I_{d})$. $U_{\alpha}(\cdot)$ is an expressive neural network, such as the UNet [Ronneberger et al., 2015]. This approach is inspired by, yet contrasts with Pang et al. [2020], wherein the latent space prior is modeled as an Energy-based Model (EBM) [Xie et al., 2016]. While EBM offers explicit unnormalized density, its sampling process is complex. Conversely, our model provides an implicit density with simpler sampling.
The trajectory generator $p_{\beta}(\tau|z)$ is a conditional autoregressive model with finite context $K$, $p_{\beta}(\tau|z)=\prod_{t=1}^{H}p_{\beta}(\tau_{(t)}|\tau_{(t-K)},...,\tau_{(t-1)},z)$ where $\tau_{(t)}=(s_{t},a_{t})$. It can be parameterized by a causal Transformer with parameter $\beta$, similar to Decision Transformer [Chen et al., 2021]. Specifically, the latent variable $z$ is included in trajectory generation using cross-attention, as shown in Fig.˜1 and controls each step of the autoregressive trajectory generation as $p_{\beta}(a_{t}|s_{t-K:t},a_{t-K:t-1},z)$. The action is assumed to follow a single-mode Gaussian distribution, i.e. $a_{t}\sim\mathcal{N}(g_{\beta}(s_{t-K:t},a_{t-K:t-1},z),I_{|A|})$.
@@ 3.2 Offline Learning
The full derivation of the learning method is in Section˜A.1. Let $\delta_{\alpha},\delta_{\beta},\delta_{\gamma}$ represent the expected gradients of $L(\theta)$ with respect to the model parameters $\alpha,\beta,\gamma$, respectively. The learning gradients for each component are formulated as follows.
@@
Estimating these expectations requires Markov Chain Monte Carlo (MCMC) sampling of the posterior distribution $p_{\theta}(z_{0}|\tau,y)$. We use the Langevin dynamics [Neal, 2011] for MCMC sampling, iterating as follows for a target distribution $\pi(z)$:
@@
The sampling process starts by initializing $z_{0}^{k=0}$ from a standard normal distribution ${\mathcal{N}}(0,I_{d})$. We then apply $N$ steps of Langevin dynamics (e.g., $N=15$) to approximate the posterior distribution, making our learning algorithm an approximate MLE. For a theoretical understanding of this noise-initialized finite-step MCMC, see Pang et al. [2020], Nijkamp et al. [2020], Xie et al. [2023]. However, for large horizons (e.g.,HH=1000), this method becomes slow and memory-intensive. To mitigate this, we adopt the persistent Markov Chain (PMC) [Tieleman, 2008, Xie et al., 2016, Han et al., 2017], which amortizes sampling across training iterations. During training, $z_{0}^{k=0}$ is initialized from the previous iteration and the number of updates is reduced to $N=2$ steps. See Section˜A.2 for training and architecture details.
@@ 3.3 Planning as Inference
This posterior sampling is achieved using Langevin dynamics similar to the training process. Specifically, we replace the target distribution in Eq.˜5 with $p_{\theta}(z_{0}|y)$ and run MCMC for a fixed number of steps. Sampling from $p_{\theta}(z_{0}|y)$ eliminates the need for expensive back-propagation through the trajectory generator $p_{\beta}(\tau|z)$.
@@
Inspired by the classifier guidance (CG) [Dhariwal and Nichol, 2021, Ho and Salimans, 2022] in conditional diffusion models, we introduce a guidance weight $w$ to the original posterior in Eq.˜6
@@
which has the score $\nabla_{z_{0}}\log\tilde{p}_{\theta}(z_{0}|y)=\nabla_{z_{0}}{\log p_{0}(z_{0})}+w\nabla_{z_{0}}{\log p_{\gamma}(y|z)}$. This guidance weight $w$ controls the interpolation between exploration and exploitation. When $w=1$, the sampled plans collectively represent the posterior density and account for Bayesian uncertainty, resulting in a provably efficient exploration scheme [Osband and Van Roy, 2017]. When $w>1$, the sampled plans are more concentrated around the modes of the posterior distribution, which are plans more likely to the agent. The larger the value of $w$, the more confident the agent becomes, and the stronger the inclination towards exploitation.
@@
Input: Learning iterations $T$, initial parameters $\theta_{0}=(\alpha_{0},\beta_{0},\gamma_{0})$, offline training samples $\mathcal{D}=\{\tau_{i},y_{i}\}_{i=1}^{n}$, posterior sampling step size $s$, the number of steps NN, and the learning rate $\eta_{0},\eta_{1},\eta_{2}$.
@@
  1.Posterior sampling: For each $(\tau_{i},y_{i})$, sample $z_{0}\sim{p}_{\theta_{t}}(z_{0}|\tau_{i},y_{i})$ using Eq.˜5 with NN steps and step-size $s$, where the target distribution $\pi$ is ${p}_{\theta_{t}}(z_{0}|\tau_{i},y_{i})$.
  2.Learn prior model $p_{\alpha}(z)$, trajectory generator $p_{\beta}(\tau|z)$ and return predictor $p_{\gamma}(y|z)$: $\alpha_{t+1}=\alpha_{t}+\eta_{0}\frac{1}{n}\sum_{i}\delta_{\alpha}(\tau_{i},y_{i})$; $\beta_{t+1}=\beta_{t}+\eta_{1}\frac{1}{n}\sum_{i}\delta_{\beta}(\tau_{i},y_{i})$; $\gamma_{t+1}=\gamma_{t}+\eta_{2}\frac{1}{n}\sum_{i}\delta_{\gamma}(\tau_{i},y_{i})$ as in Section˜3.2.
@@
Input: Expected return $y$, a trained model on offline dataset $\theta$, posterior sampling step size $s$ and the number of steps NN, Horizon $H$ and an evaluation environment.
@@
  Sample $z_{0}\sim\tilde{p}_{\theta}(z_{0}|y)$ as in Eq.˜7 using Eq.˜5 with NN steps and step size $s$, where the target distribution $\pi$ is replaced by $\tilde{p}_{\theta}(z_{0}|y)\propto p_{0}(z_{0})p_{\gamma}(y|z=U_{\alpha}(z_{0}))^{w}$ and $z=U_{\alpha}(z_{0})$.
@@
  Sample $z_{0}\sim p_{\theta}(z_{0}|y)$ as in Eq.˜6 using Eq.˜5 with NN steps and step size $s$, where $\pi$ is replaced by $p_{\theta}(z_{0}|y)\propto p_{0}(z_{0})p_{\gamma}(y|z=U_{\alpha}(z_{0}))$ and $z=U_{\alpha}(z_{0})$.
@@ 4 A Sequential Decision-Making Perspective
However, behavior cloning is believed to suffer from drifting errors since it ignores covariate shifts in future steps [Ross and Bagnell, 2010]. This concern is unique to sequential decision-making, as the agent cannot control the next state from a stochastic environment, like generating the next text token.
@@
Our modeling of the latent trajectory variable $z$ provides an alternative solution to the temporal consistency issue. Eq.˜4 is minimizing the KL divergence
@@
where $p_{\bar{\theta}}^{y}(z|\tau)={p^{y}_{\mathcal{D}}(\tau,z)}/{p^{y}_{\mathcal{D}}(\tau)}$ and $\bar{\theta}=\theta$ highlights these distributions have the same parameterization as $p_{{\theta}}^{y}$ but are wrapped with stop_grad() operator when calculating gradients for $\theta$ [Han et al., 2017]. Comparing Eqs.˜8 and 9, it is now clear that $z$ plays a similar role as $RTG$ in promoting temporal consistency in autoregressive models. Uniquely, $p_{\bar{\theta}}^{y}(z|\tau)$ is the temporal abstraction intrinsic to the model, in contrast to step-wise rewards. From a sequential decision-making perspective, $z$ is effectively a plan that the agent is persistent to. From a generative modeling perspective, $z$ from different trajectory modes would decompose the density $p^{y}(a_{t}|s_{0:t},a_{0:t-1})$, relieving the burden of learning the autoregressive policy $p_{\beta}(a_{t}|s_{0:t},a_{0:t-1},z)$.
@@
One caveat is that the transition model estimation should not be conditioned on $y$. Mixing up more trajectory regimes could provide additional regularization for its estimation and generalization. Actually, environment stochasticity is a more concerning issue for autoregressive behavior cloning, as highlighted by Yang et al. [2022], Paster et al. [2022], Štrupl et al. [2022], Brandfonbrener et al. [2022], Villaflor et al. [2022], Eysenbach et al. [2022]. Among them, Yang et al. [2022] pinpoints the issue by viewing $RTG$s as deterministic latent trajectory variables, closely related to what we present here. Uniquely, the latent variable $z$ in our model is inherently multi-modal (hence very non-deterministic) and ignorant of step-wise rewards. We postulate that the overfitting issue might be mitigated. This is validated by our empirical study inspired by Paster et al. [2022].
@@
Although RTG prediction and plan prediction both promote temporal consistency, they function very differently when mixing trajectories from multiple return-conditioned regimes. RTG prediction is a supervised learning over the joint distribution $p_{\mathcal{D}}(\tau,\rho)$. Simply mixing trajectories from multiple regimes can’t encourage generalization to trajectories that are stitched with those in the dataset. Yamagata et al. [2023] propose to resolve this by replacing $RTG$ with $RTG^{Q}$. Intuitively, this augments the distribution $p_{\mathcal{D}}(\tau,\rho)$ with $p_{\mathcal{D}}(\tau^{\prime},\rho^{Q})$, where $\tau^{\prime}$ denotes trajectories covered by the offline dynamic programming, such as Q learning, and $\rho^{Q}=(RTG_{0}^{Q},RTG_{1}^{Q},...Q_{H})$. It significantly improves tasks requiring trajectory stitching. Conversely, plan prediction is an unsupervised learning as it samples from $p_{\mathcal{D}}(\tau,y)p_{\bar{\theta}}(z|\tau,y)$. As $z$ contains more trajectory-related information than step-wise $RTG$s, trajectories lying outside of $p_{\mathcal{D}}(\tau,\rho)$ may be in-distribution for $p_{\mathcal{D}}(\tau,y)p_{\bar{\theta}}(z|\tau,y)$. The return prediction training further shapes the representation of $z$, which can be benefited from denser coverage of $y$. With more return values covered, we may count on neural networks’ strong interpolation capability to shift the trajectory distribution with $y$-conditioning.